What angle in radians is subtended by an arc 1.70m in length on the circumference of a circle of radius 2.50m ? What is this angle in degrees?
Answer #1

arc=1.7m

radius=2.5m

angle=arc/radius=0.68degrees
Answer #2

theta= s/r = 1.70 m/2.5 m = 0.68 radians(57.2957795 degrees)= 38.960